A carnot engine takes 300 calories heat from a source at 500K and rejects 150 caloies of heat to the sink,find the temperature of the sink

Grade:11

3 Answers

Samanvitha
6 years ago
In carnot engine Q1/Q2=T1/T2 HERE Q1 is 300 K Q2 IS 150K AND T1 IS 500K BY APPLYING THE FORMULA T2 IS 250 K
Varon
6 years ago
Hi christopher
apply this formula
Q1/T1 = Q2/T2
300/500 = 150/X
SO x = 250
thats it
thank you very much
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

In carnot engine Q1/Q2=T1/T2
HERE Q1 is 300 K Q2 IS 150K AND T1 IS 500K
BY APPLYING THE FORMULA
T2 IS 250 K

Thanks and Regards